Project

General

Profile

Actions

MOM update Oct 10, 2017

  • In the last week
    • 3 different cooling channel settings
      • Flip mode, 3T
      • Solenoid mode, 2T
      • Solenoid mode, 3T
      • 3-140, 6-140, 10-140 beamline settings (1M, 1M, 2M triggers at each).
      • Momentum scans 170, 200, 240 MeV
    • Field-off scattering runs overnight between ramp-downs and ramp-ups
    • Beamline tuning
  • LH2 has been very stable
  • Issues:
    • Heat load with M2D and 3T setting led to monitoring SSD pressure every 30 minutes
    • Had an overnight run with M2D in wrong polarity before it was verified & rectified
    • Rare automatic ramp-down of some BL magnets
    • 1 dead EMR SAPMT
    • Alarm handler crash once
    • Backup air compressor keeps kicking in and out -- related to when ISIS site backup compressor had issues, but that's supposedly been rectified, so not clear why our backup keeps coming on and off
    • Tracker He bubbler flow rate requires periodic check/adjustment
  • Alarm handler has been quite stable, a few more limit changes were implemented, and some more to be done
  • Plan
    • CC ramping down now for change to flip mode (with M2D)
    • Ramp up tomorrow
    • Run flip mode settings over next few days until LH2 empty next week

Updated by Rajaram, Durga about 6 years ago ยท 2 revisions